Accountancy

Assume that Sheger Manufacturing Share Company is one of the listed companies at the Addis Ababa Stock Exchange whose common stocks are traded at Br. 250 per share as at Dec. 2019. During the fiscal year of 2019, the company was able to realize net sales of Br. 50,000,000 and the Board of Directors declared, as part of their dividend policy, Br. 2,500,000 to be given to the common shareholders. There are15,000 shares outstanding common shares at Br. 150 par. Like that of last year’s performance, this year’s gross profit margin is 32% while the net profit margin is found 8%. On Dec. 31, 2019, the balance sheet of the firm reads, the current assets and the total liabilities are found 28% and 35% of the total assets respectively. Every year the share company pays interest on its long-term liabilities at a rate of 12.5% while the income tax rate of the company is 30%. Moreover, the following data are extracted from the formal financial statements of the company as of the same date.

  • Current liabilities are 45% of total liabilities

  • Total Assets turnover ratio is 2.45 times

  • Ending balances of the inventory is 35% of the current assets


Additional Information:

  • The beginning inventory is 65% of the ending inventory

  • Operating expenses (excluding interest) is 62% of gross profit amount


Assuming that there are no preferred stockholders and 365 days in a year, compute [Please round off your answers into two decimal places]


  1. Quick ratio

  2. Inventory turnover ratio in days

  3. D – E ratio

  4. Earnings Per Share

  5. Financial leverage

  6. Interest Coverage ratio

  7. Dividend Yield ratio

  8. ROA

  9. P/E ratio

  10. Market to Book Ratio
Answer #1
Total Assets turnover ratio=Sales/Total assets
Total assets=Sales/Total Assets turnover ratio=50000000/2.45=Br. 20408163
A. Quick ratio=(Current assets-Ending Inventories)/Current liabilities
Current assets=Total assets*28%=20408163*28%=Br. 5714286
Ending Inventories=Current assets*35%=5714286*35%=Br. 2000000
Current liabilities=Total liabilities*45%
Total liabilities=Total assets*35%=20408163*35%=Br. 7142857
Current liabilities=7142857*45%=Br. 3214286
Quick ratio=(5714286-2000000)/3214286=1.16
B. Inventory turnover ratio=Cost of goods sold/Average inventory
Gross profit margin=32%
Cost of goods sold=100%-32%=68%
Cost of goods sold=Total assets*68%=20408163*68%=Br. 13877551
Average inventory=(Beginning inventory+Ending inventory)/2
Beginning inventory=Ending inventory*65%=2000000*65%=Br. 1300000
Average inventory=(1300000+2000000)/2=Br. 1650000
Inventory turnover ratio=13877551/1650000=8.41
Inventory turnover ratio in days=365/Inventory turnover ratio=365/8.41=43.40 days
C. D-E ratio=Long-term liabilities/Total equity
Long-term liabilities=Total liabilities-Current liabilities=7142857-3214286=Br. 3928571
Total equity=Total assets-Total liabilities=20408163-7142857=Br. 13265306
D-E ratio=3928571/13265306=0.2962=0.30
D Earnings per share=Net profit/Common shares outstanding
Net profit=Sales*Net profit margin=20408163*8%=Br. 1632653
Earnings per share=1632653/15000=Br. 108.84
E Financial leverage=Total liabilities/Total equity=7142857/13265306=0.54
F Interest Coverage ratio=Operating expenses/Interest expense
Operating expenses=Gross profit*62%
Gross profit=Sales*Gross profit margin=20408163*32%=$ 6530612
Operating expenses=6530612*62%=$ 4048979
Interest expense=Long-term liabilities*Interest rate=3928571*12.5%=Br. 491071.4
Interest Coverage ratio=4048979/491071.4=8.25 times
G. Dividend yield=Dividends per share of common stock/Market price per share
Dividends per share of common stock=Dividends on common stock/Number of common stock outstanding=2500000/15000=Br. 166.67
Dividend yield=166.67/250=0.66668=66.67%
H. ROA=Net profit/Total assets=1632653/20408163=0.08
I. P/E ratio=Market price per share/Earnings per share=250/108.84=2.30
J. Market to Book Ratio=Market price per share/Net book value per share
Net book value per share=Total equity/Number of common stock outstanding=13265306/15000=Br. 884.35
Market to Book Ratio=250/884.35=0.28
Answer #2

A. QUICK RATIO = (Current assets - inventory) \div Current liabilities   Current assets = 28% of Total assets Total assets = Net sales \div Total assets turnover ratio = 50,000,000 \div 2.45 = 20,408,163 Current assets = 28% of 20,408,163 = 5,714,285 Inventory = 35% of current assets = 35% of 5,714,285 = 2,000,000 Current liabilities = 45% of Total liabilities Total liabilities = 35% of Total assets So, current liabilities = 45% of 35% of Total assets    = 45% of  35% of 20,408,163 = 3,214,285 Quick ratio = (5,714,285 - 2,000,000) \div 3,214,285 = 3,714,285 \div 3,214,285 = 1.16 (approx)

B. INVENTORY TURNOVER RATIO in days = Average Inventory \times 365 \div Cost of goods sold Average inventory = (opening inventory + ending inventory) / 2 Opening inventory = 65%of ending inventory = 65% of 2,000,000 = 1,300,000 Average inventory = (1,300,000 + 2,000,000) / 2 = 1,650,000 Cost of goods sold = Sales - Gross profit = 50,000,000 - 32% of 50,000,000 = 50,000,000 - 16,000,000 = 34,000,000 Inventory turnover ratio in days = 1,300,000 \times 365 \div 34,000,000 = 13.95 days (approx)